after a 20% discount ,you purchase a television for 240.what wasthe televisions price before the reduction?define a variable,create an equation, solve using algebra,and answers in a sentence
Answer by Theo(13342)

x equal the original price.

x - .2 * x = .8 * x equals the discount price.

you buy the television for the discount price.

therefore .8 * x = 240.

solve for x to get x = 240 / .8 = 300.

the original price is 300.

20% discount off of 300 is equal to 60.

discount price is equal to 300 - 60 = 240.

solution is:

the original prices before the 20% discount was 300 dollars.
the discount of 20% was applied to 300 dollars, resulting in a reduction in the original price of 60 dollars.
this resulted in a discounted price of 240 dollars.